OpenStudy (anonymous):

im buying a home for 50.000. with 9% interest .i pay 450.a month how many years will it take to pay off

OpenStudy (anonymous):

50,000*1.09-450x=0 54500-450x=0 +450x +450x 54500=450x divide both sides by 450 121.11111=x